Ticket Number: HEA-NOC/20000908-01 Ticket Status: CLOSE
Ticket Type: Unscheduled Ticket Source: HEA-NOC
Ticket Scope: Site Site/Line: Andel
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: Problem Start: 20000908 10:49 UTC+1
Ticket Closed: 20000920 09:18:43 UTC Problem End: 20000920 10.10 UTC+1
We have shutdown the London VC on our router Andel, as it is pending
All traffic is re-routing via the Paris VC for TEN-155.
All traffic is re-routing via the Paris VC, so there have been no outages.
20000920 10.10 UTC+1
The memory upgrade on andel has been completed and this line has been restored. Ticket closed.
Time to Fix:
11 days, 23hrs and 39 mins. There were no outages during this time as all
traffic was backed up via our Paris VC.